The solubility of adipic acid in water is 1.40 g/ 100 mL at 10 C, and 160 g/ 100 mL at 100 C.

A. If 0.950 g of adipic acid is dissolved in 3.00 mL of boiling water and cooled to 10 C. Calculate the maximum mass of adipic acid that could crystallize from the solution.

B. What is the maximum amount of adipic acid that could dissolve in 3.00 mL of boiling water?
Answer #1

A) Solubility is the mass of solute that can be dissolved in 100g of the solvent at a particular temperature (mass= volume, for water as density is 1g/ml)

ie Solubility = (Mass of solute /Mass of solvent ) x 100

Given solubility at 10 0C = 1.40g/100 ml

Solubility at 100 0C = 160 g/100ml

Given solution has 0.950 g adipic acid dissolved in 3 ml water at 100 0C. It is cooled to 100C

Let 'x' be the mass of adipic acid that can be dissolved in 3 ml water at 100 C

Then 1.40 g = (x/3 ml ) x 100 ml

x = (1.40g x 3 ml )/100 ml =0.042 g

Mass of adipic acid that could be crystallised = 0.950 g - 0.042 g = 0.908 g

B) Let 'y' be the maximum amount of adipic acid that can be dissolved in 3 ml water at 100 0 C

(y/3ml) x 100 ml = 160 g

y = (3ml x 160 g ) /100 ml = 4.8 g

Mass of adipic acid = 4.8 g